10 kg of air is contained in a constant volume container at a temperature of AA°C and a pressure of 150kPa. How much air is pressed into the container, its pressure is 250kPa and its temperature is BB°C? (AA and BB are a two-digit number ending in 0 or 5)


1
Expert's answer
2021-12-27T07:55:33-0500

"p_1V=\frac mMRT_1,"

"p_2V=\frac{ m+dm}MRT_2,\implies"

"\frac{p_2}{p_1}=(1+\frac{dm}m)\frac{T_2}{T_1},"

"dm=m(\frac{p_2T_1}{p_1T_2}-1)."